OpenStudy (anonymous):

Which of the ordered pairs in the form (x, y) is a solution of this equation? 6x-y/2=14 (-9, 3) (2, -4) A. Neither is a solution. B. The first is a solution, but the second is not. C. Both are solutions. D. The first is not a solution, but the second is.

OpenStudy (anonymous):

try plugging the points in the equation and whichever point satisfies the equation that will be the right answer

OpenStudy (campbell_st):

just substitute each pair and see if they are true (-9, 3) so x = -9 and y = 3 \[6 \times (-9) - \frac{3}{2} = -54 - \frac{3}{2} = ?\] will this equal 14...? next point (2, -4) x = 2 and y = -4 \[6 \times 2 - \frac{(-4)}{2} = 12 + \frac{4}{2} = ?\] which ordered pair has 14 as an answer

OpenStudy (anonymous):

the first one?